# Augmented Neural Fine-Tuning for Efficient Backdoor Purification

### v1

@@ 1 Introduction
Machine learning and computer vision algorithms are increasingly common in safety-critical applications [20, 3], necessitating the design of secure and robust learning algorithms. Backdoor attack [9, 17] on deep neural network (DNN) models is one of the heavily studied branches of AI safety and robustness. Backdoor defenses can generally be categorized into two major groups based on whether the defense is done during training or test. Training time defense (e.g.[27, 49, 25]) focuses on training a benign model on the poisonous data, while test time defense (e.g.[34, 54, 6, 29]) deals with purifying backdoor model after it has already been trained. In this work, our goal is to develop an efficient test-time defense. Some test-time defenses focus on synthesizing trigger patterns [51, 7, 52] followed by vanilla weight fine-tuning. Most of these defenses aim to synthesize class-specific triggers independently or use additional models to generate the triggers simultaneously. Recent state-of-the-art (SOTA) backdoor defense methods, e.g., ANP [54], I-BAU [59], AWM [6], employ very similar techniques and do not work well without an expensive adversarial search module. For example, ANP [54] performs an adversarial search to find vulnerable neurons responsible for backdoor behavior. Identifying and pruning vulnerable neurons require an exhaustive adversarial search, resulting in high computational costs. Similar to ANP, AWM and I-BAU also resort to trigger synthesizing with a modified adversarial search process. Another very recent technique, FT-SAM [66] uses sharpness-aware minimization (SAM) [15] to fine-tune model weights. SAM is a recently proposed optimizer that utilizes Stochastic Gradient Descent (SGD), which penalizes sudden changes in the loss surface by constraining the search area to a compact region. Since SAM performs a double forward pass to compute the loss gradient twice, it results in a notable runtime increase for FT-SAM. In our work, we aim to develop an effective backdoor defense system that neither requires an expensive adversarial search process to recover the trigger nor a special type of optimizer with a runtime bottleneck.
@@ 3 Threat Model
$\mathcal{L}({\theta^{*}},\mathbb{D}_{\mathrm{train}})=\frac{1}{N}\sum_{i=1}^{N}[\ell(y_{i},f_{\theta^{*}}(x_{i}))],$ (1)
@@ 4.2 Clean Accuracy Retainer
$\widehat{M}=\operatorname*{arg\,min}_{M~{}\mid~{}m_{l}^{(i)}\in[\mu(l),1],~{}\forall l,i}\mathcal{L}^{\mathrm{mix}}({\theta\odot M},\mathbb{D}_{\mathrm{val}})$ (6)
@@
where $\theta:=\{\Theta_{1},\ldots,\Theta_{L}\}$, $M:=\{M_{1},\ldots,M_{L}\}$, and $M_{l}=[m_{l}^{(1)}\cdots m_{l}^{k_{l}}]^{T}\in\mathbb{R}^{k_{l}}$. We have

### v2

@@ 1 Introduction
Machine learning and computer vision algorithms are increasingly common in safety-critical applications [27, 4], necessitating the design of secure and robust learning algorithms. Backdoor attack [10, 22] on deep neural network (DNN) models is one of the heavily studied branches of AI safety and robustness. Backdoor defenses can generally be categorized into two major groups based on whether the defense is done during training or test. Training-time-defense (e.g.[34, 59, 32]) focuses on training a benign model on the poisonous data, while test-time-defense (e.g.[41, 65, 7, 36]) deals with purifying backdoor model after it has already been trained. In this work, our goal is to develop an efficient test-time defense. Some test-time defenses focus on synthesizing trigger patterns [62, 8, 63] followed by vanilla weight fine-tuning. Most of these defenses aim to synthesize class-specific triggers independently or use additional models to generate the triggers simultaneously. Recent state-of-the-art (SOTA) backdoor defense methods, e.g., ANP [65], I-BAU [71], AWM [7], employ very similar techniques and do not work well without an expensive adversarial search module. For example, ANP [65] performs an adversarial search to find vulnerable neurons responsible for backdoor behavior. Identifying and pruning vulnerable neurons require an exhaustive adversarial search, resulting in high computational costs. Similar to ANP, AWM and I-BAU also resort to trigger synthesizing with a modified adversarial search process. Another very recent technique, FT-SAM [78] uses sharpness-aware minimization (SAM) [19] to fine-tune model weights. SAM is a recently proposed optimizer that utilizes Stochastic Gradient Descent (SGD), which penalizes sudden changes in the loss surface by constraining the search area to a compact region. Since SAM performs a double forward pass to compute the loss gradient twice, it results in a notable runtime increase for FT-SAM. In our work, we aim to develop an effective backdoor defense system that neither requires an expensive adversarial search process to recover the trigger nor a special type of optimizer with a runtime bottleneck.
@@ 3 Threat Model
$\mathcal{L}({\theta^{*}},\mathbb{D}_{\mathrm{train}})=\frac{1}{N}\sum_{i=1}^{N}[\ell(y_{i},f_{\theta^{*}}(x_{i}))].$ (1)
@@ 4.2 Clean Accuracy Retainer
$\widehat{M}=\operatorname*{arg\,min}_{M~{}\mid~{}m_{l}^{(i)}\in[\mu(l),1],~{}\forall l,i}\mathcal{L}^{\mathrm{mix}}({\theta\odot M},\mathbb{D}_{\mathrm{val}}).$ (6)
@@
where $\theta:=\{\Theta_{1},\ldots,\Theta_{L}\}$, $M:=\{M_{1},\ldots,M_{L}\}$, and $M_{l}=[m_{l}^{(1)}\cdots m_{l}^{(k_{l})}]^{T}\in\mathbb{R}^{k_{l}}$. We have